OpenStudy (anonymous):

use the integral test to determine whether the infinite series sum n=1 to infinity of n(e)^-n^2

OpenStudy (anonymous):

I've still got some time left, so I can give you a short hint. When integrating, let \(u=n^2\).

OpenStudy (anonymous):

and -du=2n

OpenStudy (anonymous):

please guys I really need help with this problem its driving me crazy

OpenStudy (anonymous):

\[\int_1^\infty \frac{n}{e^{2n}}~dn\] If the integral converges/diverges, then so does the series. Letting \(u=n^2~\Rightarrow~du=2n~dn\), so \[\frac{1}{2}\int_1^\infty e^{-u}~du\] Does the integral converge or diverge?

OpenStudy (anonymous):

\[\bf= \frac{ 1 }{ 2 }\lim_{n \rightarrow \infty}\int\limits_{1}^{n}e^{-u} \ du=\frac{ 1 }{ 2 }\lim_{n \rightarrow \infty}-e^{-n}+e^{-1}\]Can we evalute this limit? @Brittny

OpenStudy (anonymous):

@Brittny ???

OpenStudy (anonymous):

im sorry just gimme a sec

OpenStudy (anonymous):

im sloving it

OpenStudy (anonymous):

it think it will be 1/2 -e^(- infinity)+e-1

OpenStudy (anonymous):

it divergent because it goes to infinty , right?

OpenStudy (anonymous):

@genius12 is it right ?

OpenStudy (anonymous):

@Brittny I think it is best if you look at these two places here: http://www.youtube.com/watch?v=ojztrQMqLgE http://cims.nyu.edu/~kiryl/Calculus/Section_8.3--The_Integral_and_Comparison_Tests/The_Integral_and_Comparison_Tests.pdf They're both extremely helpful.

OpenStudy (anonymous):

ok after looking at the problem again. I solved it and got -1/4e^4 which is by the p test converges right ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

@Brittny I suggest you watch the video through the link; the guy does this exact same problem.